4. (2 points) Solve these recurrence relations together with the initial conditions given. Each is worth 1 point.

a) an = 6an-1 - 11an-3 +6an-3 for n > 3, a0 = 4, a1 = 6, a2 = 12

b) an = an–1 +9an-2 – 9an-3 for n > 3, a0 = 6, a1= 0, a2 = 30


1
Expert's answer
2021-11-22T10:49:31-0500

a)

The characteristic polynomial of this recurrence relation is


"r^3-6r^2+11r-6=0"

"r^2(r-1)-5r(r-1)+6(r-1)=0"

"(r-1)(r^2-5r+6)=0"

"(r-1)(r-2)(r-3)=0"

"r_1=1, r_2=2, r_3=3"

Hence, the solutions to this recurrence relation are of the form


"a_n=\alpha_1\cdot1^n+\alpha_2\cdot2^n+\alpha_3\cdot3^n"

Use the initial conditions


"a_0=4=\alpha_1+\alpha_2+\alpha_3""a_1=6=\alpha_1+2\alpha_2+3\alpha_3""a_2=12=\alpha_1+4\alpha_2+9\alpha_3"

"\alpha_1+\alpha_2+\alpha_3=4""\alpha_2+2\alpha_3=2""2\alpha_2+6\alpha_3=6"

"\alpha_1+\alpha_2+\alpha_3=4""\alpha_2+2\alpha_3=2""\alpha_2+3\alpha_3=3"


"\alpha_1=3""\alpha_2=0""\alpha_3=1"

Hence, the unique solution to this recurrence relation and the given initial conditions is the sequence "\{a_n\}" with


"a_n=3+3^n"

b)

The characteristic polynomial of this recurrence relation is


"r^3-r^2-9r+9=0"

"r^2(r-1)-9(r-1)=0"

"(r-1)(r-3)(r+3)=0"

"r_1=1, r_2=-3, r_3=3"

Hence, the solutions to this recurrence relation are of the form


"a_n=\alpha_1\cdot1^n+\alpha_2\cdot(-3)^n+\alpha_3\cdot3^n"

Use the initial conditions


"a_0=6=\alpha_1+\alpha_2+\alpha_3""a_1=0=\alpha_1-3\alpha_2+3\alpha_3""a_2=30=\alpha_1+9\alpha_2+9\alpha_3"

"\alpha_1+\alpha_2+\alpha_3=6""4\alpha_2-2\alpha_3=6""8\alpha_2+8\alpha_3=24"

"\alpha_1+\alpha_2+\alpha_3=6""2\alpha_2-\alpha_3=3""\alpha_2+\alpha_3=3"


"\alpha_1=3""\alpha_2=2""\alpha_3=1"

Hence, the unique solution to this recurrence relation and the given initial conditions is the sequence "\{a_n\}" with


"a_n=3+2\cdot(-3)^n+3^n"